OpenStudy (anonymous):

I got the answer solved for S1, but am having problems with S2 adn S3, making them true. I am following all the steps but it just won't work! Sn: 1^2+4^2+7^2+...+(3n-2)^2 = n(6n^2-3n-1)/2.

OpenStudy (phi):

what did you get for S1 ?

OpenStudy (anonymous):

1=1

OpenStudy (anonymous):

I just plugged in 1^2

OpenStudy (phi):

you should "plug in" 1 (not 1^2) (though 1^2 = 1 , so it still works)

OpenStudy (anonymous):

Let me try that first. Hold on.

OpenStudy (phi):

S2 is short for use n=2 in the formula n(6n^2-3n-1)/2

OpenStudy (anonymous):

Okay for S2, I got 1,1) so it is true.2

OpenStudy (anonymous):

So for S3, I assume I plug in 2?

OpenStudy (phi):

S1 means use n=1 (and you will get 1 from the formula) for S2, use n=2 what do you get ?

OpenStudy (anonymous):

Hold on. For S1, I got 1, 1. Let me try 2.

OpenStudy (anonymous):

using 2 in S2, I got 16=18. Not true.

OpenStudy (anonymous):

Sorry, I meant 16=17

OpenStudy (phi):

the "left side" goes from 1^2 upto (3n-2)^2 when n=1 the last term is (3*1-2)^2 = (3-2)^2 = 1^2 so the series for S1 is just S1= 1^2 for S2, the last term is (3*2 - 2)^2 = (6-2)^2 = 4^2 and the series is S2= 1^2 + 4^2 if we simplify that we get 1+16= 17 so S2 = 17 now use the formula (the right side) replace n with 2 and get \[ \frac{2(6\cdot 2^2-3\cdot 2-1)}{2} \]

OpenStudy (anonymous):

Having battery issues with calculator. Hold on.

OpenStudy (anonymous):

16=16

OpenStudy (phi):

no, you should get 17 remember, the series is a sum of n terms in other words, when n=2 , S2 is the sum of 2 terms: 1^2 + 4^2 which adds up to 17 (1+16) the formula, if you are careful, also gives 17

OpenStudy (phi):

do it step by step. for n=2 the formula (after replacing all the n's with 2, becomes \[ \frac{2(6\cdot 2^2-3\cdot 2-1)}{2} \] inside the parens we have 6*2*2 - 6 - 1 or 24 -7 or 17 so you have \[ \frac{2(17)}{2} \]

OpenStudy (anonymous):

Okay. that is interesting.

OpenStudy (anonymous):

I just redid it and got 17=17. Thanks.

OpenStudy (anonymous):

Thanks.

OpenStudy (phi):

can you do S3? the left side will have 3 terms 1^2+ 4^2 + 7^2

OpenStudy (phi):

fyi, when they write Sn: 1^2+4^2+7^2+...+(3n-2)^2 the (3n-2)^2 is the *last term* when n is 3, we get the last term= (3*3-2)^2 = (9-2)^2 = 7^2 that means S3 is S3= 1^2+4^2+7^2 also, S4= 1^2+4^2+7^2+10^2, and S5= 1^2+4^2+7^2+10^2 + 13^2 and so on.

OpenStudy (anonymous):

Thank you. I just did S3 and got 67=67.

OpenStudy (phi):

I get 66

OpenStudy (phi):

the left side is S3= 1^2+4^2+7^2 = 1 + 16 + 49 = 50+16 = 66

OpenStudy (anonymous):

Let me retry. I think I must of miscalculated a number someplace.

OpenStudy (phi):

it helps to write it down, and simplify in small steps.

OpenStudy (anonymous):

I still get 67=67. I will continue trying. Thank you phi. I do appreciate your help. But I will keep trying. Let me take a break and retry it again. Sometimes you need to walk away for a little bit. Been at 3 problems for 3 hours and this is the one wiht the problem. Thanks.

OpenStudy (phi):

how do you get 67 from the left side which is 1^2+4^2+7^2 ?